OpenStudy (anonymous):

0.5(y+2)-0.3=0.4+0.5 My book says the answer is -2 but I keep coming up with the same 2 different answers depending on the way I work it

OpenStudy (anonymous):

I think you may have made a typo in typing the equation, because .4 + .5 does not look right. The way it is written there, y = .4.

OpenStudy (anonymous):

let's see. 0.5y+1-0.3=0.9 (here i distributed, then added) 0.5y+0.7=0.9 0.5y=0.2 hm... I get 2/5

OpenStudy (anonymous):

i agree... check the problem. i got the same answer (2/5=0.4)

OpenStudy (anonymous):

I think he meant = .4y + .5. That produces an answer of -2.

OpenStudy (anonymous):

what i wrote is what the book says...is the answer they gave wrong because i have also gotten 2.5

OpenStudy (anonymous):

Anyway, explanation, .5(y+2) -.3 = .4y + .5 .5y + 1 - .3 = .4y + .5 (Distribution) .5y + .7 = .4y + .5 .5y = .4y - .2 .1y = -.2 y = -.2/.1 y = -2 If that's what your book says, I think the typo is in the book. Point it out to your teacher for extra credit.

OpenStudy (anonymous):

0.5 is the same as .5. so 1-0.3 is the same as 1.3 which is 0.7. I double checked on my calculator

OpenStudy (anonymous):

either way the answer is the same. thanks for showing me where i went wrong. that was a huge relief. lol